OpenStudy (destinyyyy):

Can someone please help me with these last 5 problems? Ive been working on this all day.

OpenStudy (destinyyyy):

Solve using appropriate substitution. (x^2 -6x)^2 -11(x^2 -6x) -30=0

OpenStudy (destinyyyy):

Im at the end where I factor for the second time

OpenStudy (destinyyyy):

x^2 -6x-5=0 x^2 -6x-6=0

OpenStudy (destinyyyy):

(x-1)(x-5)=0 I cant seem to factor the second one.. I dont get why

OpenStudy (tkhunny):

Have you met the "Discriminant"? \((-6)^{2} - 4(1)(-5) = 36 + 20 = 56\) 56 is not a perfect square. You will not be able to factor it. Your factorization is not correct. \((-6)^{2} - 4(1)(-6) = 36 + 24 = 60\) 60 is not a perfect square. You will not be able to factor it.

OpenStudy (destinyyyy):

????

OpenStudy (destinyyyy):

Where are you getting all that from?

OpenStudy (destinyyyy):

@freckles

OpenStudy (tkhunny):

You didn't answer my first question. If you have not met the "Discriminant", that will not make much sense. Anyway, (x-1)(x-5) = x^2 - 6x + 5 -- The sign on the constant term is not correct as you have it factored. Have you encountered ANY situation where things cannot be factored? There are other ways to proceed when factoring fails.

OpenStudy (destinyyyy):

I have in the last section the other day..

OpenStudy (destinyyyy):

Yes but all my previous examples and help from people have only factored so I didnt think I was allowed to do a different way

OpenStudy (tkhunny):

Well, we're stuck on these. They can't be factored. I hope the assignment was not meant just to frustrate you. It's time to pull out the Quadratic Formula or "Completing the Square" or anything else you have in your experience.

OpenStudy (destinyyyy):

Okay

OpenStudy (destinyyyy):

3+- square root 15

OpenStudy (destinyyyy):

3+- square root 14

OpenStudy (destinyyyy):

My answers are wrong.

OpenStudy (destinyyyy):

@IrishBoy123

OpenStudy (tkhunny):

Since you haven't told us what the question is, it's pretty tough to help you with that. You also haven't shown your work.